If a force acts on an object only to accelerate it what is true about the equations: w=Fs and Ek =1/2mv^2-1/2mu^2
1
Expert's answer
2019-05-27T11:36:08-0400
"W=Fs"

is true only if F is the constant force.


"\Delta E_k =\frac{1}{2}mv^2-\frac{1}{2}mu^2"

is the formula for the change in the kinetic energy.

We have:


"\Delta E_k =W"

"\frac{1}{2}mv^2-\frac{1}{2}mu^2=Fs"